Lucy is four times as old as Mary. In twenty years Lucy will be three times as old as Mary. How old are they now?
Answer by sofiyac(983)

Mary is x
Lucy is 4x
3(x+20)=4x+20
3x+60=4x+20
40=x, then lucy is
4*40=160
so mary is 40, lucy is 160, in twenty years they'll be 60 and 180, so lucy would be 3 times as old as mary. Although i dont know of anyone that lived that long. So the answer is mary is 40 and lucy is 160